three years ago the sum of the ages of father and his son was 48 years hence from now father's age will be three times of his son. find the present age of his son.
Found 2 solutions by Boreal, addingup:
Answer by Boreal(15235)

son's age now=x
son's age 3 years ago= x-3
father's age 3 years ago was such that x-3+f=48;f=51-x
Therefore, father's age now is 51-x+3=54-x
I am assuming now the father's age is 3 times that of his son.
54-x=3x
4x=54
x=13.5 years ANSWER
father's age now is 40.5 years.
3 years ago, the ages were 10.5 and 37.5, and they add to 48.

Answer by addingup(3677)

"hence from now father's age will be three times of his son" How many years from now, please?